Calculate the rate constant at 225 degrees C for a reaction that has a rate constant of 8.1 x 10^-4 s^-1 at 95 degrees C and an activation energy of 95.0 kj/mol
Answer #1

Ahrrenius equation

ln(k2/k1)= E/R*(1/T1-1/T2)

ln(k2/(8.1*10^-4)) = (95000/8.314)*(1/(95+273) - 1/(225+273))

k2 =  (8.1*10^-4)*exp((95000/8.314)*(1/(95+273) - 1/(225+273)))

K2 = 2.683234